A man wanted to measure the height of a building by dropping a bob from the top. If the bob hits the ground 2.30 seconds after it was dropped, how high is the building? 



1
Expert's answer
2021-11-11T11:59:28-0500

The height is


"h=\frac{gt^2}2=\frac{9.8·2.3^2}{2}=25.9\text{ m}."